# Parametrized multiqubit gate design for neutral-atom based quantum platforms

### v1

@@ Abstract
A clever choice and design of gate sets can reduce the depth of a quantum circuit, and can improve the quality of the solution one obtains from a quantum algorithm. This is especially important for near-term quantum computers that suffer from various sources of error that propagate with the circuit depth. Parametrized gates in particular have found use in both near-term algorithms and circuit compilation. The one- and two-qubit versions of these gates have been demonstrated on various computing architectures. The neutral atom platform has the capability to implement native $N$-qubit gates (for $N\geq 2$). However, one needs to first find the control functions that implement these gates on the hardware. We study the numerical optimization of neural networks towards obtaining families of controls – laser pulses to excite an atom to Rydberg states – that implement phase gates with one and two controls, the $\mathrm{C_{1}P}$ and $\mathrm{C_{2}P}$ gates respectively, on neutral atom hardware. The pulses we obtain have a duration significantly shorter than the loss time scale, set by decay from the Rydberg state. Further, they do not require single-site addressability and are smooth. Hence, we expect our gates to have immediate benefits for quantum algorithms implemented on current neutral atom hardware.
@@ I Introduction
To implement these gates on hardware, however, one must devise pulses – that is, the time profile of the laser intensity and frequency – that can carry out the desired unitary transformation on the electronic states of the atoms. There exist both theoretical proposals, notably for $k$-qubit Toffoli and fan-out gates (for $k>2$) [13, 14, 15, 16, 6], as well as recent experimental demonstrations of these gates for the Rydberg platform [15, 17, 18]. Notably, Cao et al. implemented $N$-qubit rotation gates (for $N$ up to 9) in order to generate cascaded GHZ states [17] by placing $N$ atoms within interacting range of each other. The pulses devised to implement this gate are also time-optimal – meaning they are the shortest possible pulses that implement the corresponding unitary. Such pulses that implement the gate directly are significantly shorter in duration than implementing the decomposition of the gate in terms of one- and two-qubit gates [19], and are obtained numerically using a class of methods termed quantum optimal control (QOC) [13, 20].
@@
Recent studies have also investigated the parametrized versions of such multiqubit gates – gates that implement some unitary $U(\bm{\phi})$, for $M$ parameters $\bm{\phi}=(\phi_{1},\phi_{2},\dots,\phi_{M})$. In terms of variational algorithms, there are proposed applications of such gates towards simulating lattice gauge theories [21] and combinatorial optimization [22]. Further, for a given circuit, the inclusion of $k$-controlled phase, or $\mathrm{C}_{k}\mathrm{P}$ gates (for $k\geq$ 2) in the gate set can reduce the gate count and the total circuit evaluation time of the compiled circuit [12], thus reducing infidelities – as most error sources propagate in time [13]. This reduction is especially significant for the current era of quantum hardware, termed the noisy intermediate-scale quantum (NISQ) era. In these devices, the coherence times of the qubits are limited, and there are multiple sources of gate infidelity.
We investigate the problem of finding families of pulses that implement the $\mathrm{C_{1}P}$ and $\mathrm{C_{2}P}$ gates on Rydberg hardware. Employing numerical techniques from quantum optimal control, we find pulses for both gates that are fast – more specifically, the $\mathrm{C_{1}P}$ pulse we obtain is time-optimal, and the $\mathrm{C_{2}P}$ pulse operates on a time scale comparable to the time-optimal solution. These pulses are also smooth, and require only global addressing – the same pulse is sent to all atoms – and hence can be implemented on NISQ devices.
We note that although preliminary work exists on $\mathrm{C_{1}P}$ gate protocols for the Rydberg platform (which we will discuss in Sec. II.1), the $\mathrm{C_{2}P}$ protocol has not been explored to the best of our knowledge. Hence, our paper presents the first $\mathrm{C_{2}P}$ protocol for the Rydberg platform. Further, the techniques we use for the optimization can be extended to gates over more control qubits, as well as to fast quantum state preparation for parametrized states.
@@ II.1 Multiqubit gates
One can always implement such transformations by decomposing them into a product of one- and two-qubit gates available on the hardware – this is visualized for the $\mathrm{C_{1}P}$ gate in Fig. 2. However, this approach has a significantly longer gate duration [19, 26] than the control-level approach we adopt. In Appendix C, we present an in-depth comparison of these two approaches for our problem.
We note that as $\mathrm{P(\pi)=Z}$, $\mathrm{C}_{k}\mathrm{P}(\pi)$ is equivalent to a $\mathrm{C}_{k}\mathrm{Z}$ gate. Hence, the problem we consider is the generalized version of the problem of obtaining a $\mathrm{C}_{k}\mathrm{Z}$ pulse [13, 27, 28, 29].
@@
and
$H_{\mathrm{int}}^{i,j}=-\frac{C_{6}}{|\bm{R}_{i}-\bm{R}_{j}|^{6}}(\mathds{1}_{1}\otimes\mathds{1}_{2}\otimes\dots\otimes\ket{r}\bra{r}_{i}\otimes\dots\ket{r}\bra{r}_{j}\otimes\dots\mathds{1}_{N}).$ (10)
@@
For an $N$-qubit system in this regime that evolves under the Hamiltonian from Eq. (8), one can realize unitaries that map each computational basis state to itself up to a phase [29], which makes it well suited to realizing $\mathrm{C}_{N-1}\mathrm{P}$ gates [12] – upto a global single qubit $\mathrm{R_{Z}}$ rotation. The dynamics in the blockade regime can be understood in terms of two-level systems (TLS): two and three for the $\mathrm{C_{1}P}$ and $\mathrm{C_{2}P}$ gate, respectively. We illustrate these systems in Fig. 1. The TLS are closed systems – independent of each other, they partition the Hilbert space into subspaces with populations that do not mix [29].
@@
We note, however, that the blockade strength is finite in the experiment [as computed in Eq. (10)], and this is a potential source of infidelity for gates optimized at $B\to\infty$. For the $\mathrm{C_{1}P}$ gate, we will hence simulate the dynamics (during the optimization) under the full Hamiltonian from Eq. (8) with $B=21.1$ [27], an experimentally realistic value – as opposed to simulating $H_{\mathrm{eff}}$. This is done in order to mitigate the finite blockade errors [28].
@@ II.2 Numerical optimization
Sauvage et al. investigated the generation of parametrized gates by training neural networks (NNs) – used as variational ansatze – to model families of pulses [19]. We build upon their approach, and adapt it to our neutral atom system. As we will motivate in this section, compared to GRAPE, the NN only has to be trained once, and then outputs a high-fidelity pulse for any angle in the domain of the gate without requiring further optimizations.
@@
for $\mathcal{T}$ the time-ordering operator. The cost function at each iteration is
$J=1-\frac{1}{4^{(k+1)}}\braket{}{U_{\mathrm{C}_{k}\mathrm{P}}(\phi)^{{\dagger}}(PU_{\mathrm{out},k}(\phi)P)}{{}^{2}}_{\phi}$ (15)
where $P$ is the projection operator onto the computational subspace [35], ||.||||.||| | . | | the Hilbert-Schmidt norm and .$\braket{.}_{\phi}$ implies we average over samples of $\phi$ drawn randomly from $\mathrm{dom}(\mathrm{C}_{k}\mathrm{P})$ in each iteration. We will refer to $J=\braket{(1-F)}_{\phi}$ as the average infidelity over the domain of the protocol. We then train the weights of the NN to minimize the cost function.
@@ III Results
To compute the infidelity of our pulses in a realistic experimental setup, we simulate decay from the Rydberg state, with estimated lifetime $\tau=96.5~{}\mathrm{\mu s}(\simeq 6\times 10^{3}/\Omega_{\mathrm{max}})$ for the $n=61$ state [28], considering a Rabi frequency of $\Omega_{\mathrm{max}}=2\pi\times 10\mathrm{MHz}$ for Sr, which is within reach in the near term [3, 27].
As discussed in Sec. II.1, we also consider a realistic blockade strength of $B=21.1$ [27], unless specified otherwise. Calculations of infidelities arising from the decay of |$\ket{r}$, and from finite blockade, are discussed in Appendix B. The total mean infidelity $\braket{(1-F)}_{\phi}$ was computed over 100 samples randomly drawn from $\mathrm{dom}(\mathrm{C}_{k}\mathrm{P})$.
@@ III.1 $\mathrm{C_{1}P}$ gate
Simulating the errors as discussed above, we obtain a mean infidelity $\braket{(1-F)}_{\phi}=3.4\times 10^{-4}$ for the two-qubit $\mathrm{C_{1}P}$ gate over $\mathrm{dom(C_{1}P)}$ – with the main error arising due to Rydberg state decay (see inset of Fig. 4(b)).
@@ III.2 $\mathrm{C_{2}P}$ gate
We hence study the strategy of using the $\mathrm{C_{2}P}$ NNs we trained at $B\to\infty$, as initial guesses for $\mathrm{C_{2}P}$ NNs trained at $B=21.1$, and then re-run the optimization as discussed in Sec. II.2. The pulses we obtain from this optimization reduce the average finite blockade infidelity to $\braket{(1-F)_{\mathrm{int}}}_{\phi}=4\times 10^{-4}$, and the total infidelity $\braket{(1-F)}_{\phi}=1.2\times 10^{-3}$ – where the decay from the Rydberg state is the main source of infidelity. The results of this optimization are visualized in Fig. 6, in Appendix B.
@@ Appendix B Losses from finite blockade and Rydberg state decay
To simulate the effects of spontaneous emission and blackbody radiation (BBR) at $300\mathrm{K}$ from the Rydberg state, we add a non-Hermitian term [27, 28]
$H_{\mathrm{decay}}=i\frac{\Gamma_{n=61}}{2}\ket{r}\bra{r},$ (21)
to $H_{1q}$ in Eq. (7), which assumes that all decay occurs to states outside of the computational subspace – an assumption justified for the system we studied in this article [27, 52]. This term is subsequently used to construct the full Hamiltonian $H_{Nq,\Gamma}$, in a manner similar to Eq. (8). Here $\Gamma_{n=61}$ is the decay width of the $5\mathrm{s}61\mathrm{s}\,^{3}\mathrm{S}_{1}$ state in Sr, corresponding to a lifetime $1/\Gamma_{n=61}=96.5\mathrm{\mu s}$ [28].
@@
for $F,F_{\mathrm{decay}}$ the fidelities are obtained by simulating the pulses under $H_{Nq}$ [see Eq. (8)], and $H_{Nq,\Gamma}$ respectively. The infidelity arising from finite blockade was defined for three qubits, in a manner similar to Eq. (22), $(1-F)_{\mathrm{int}}=F_{\infty}-F_{\mathrm{fin}}$, for $F_{\infty},F_{\mathrm{fin}}$ the fidelities of the pulses simulated for $H_{\mathrm{eff}}$ and $H_{3q}$ respectively. This infidelity is visualized as a function of $\phi$ for in Fig. 6.
@@
Appendix C Decomposition and native gates

### v2

@@ Abstract
A clever choice and design of gate sets can reduce the depth of a quantum circuit, and can improve the quality of the solution one obtains from a quantum algorithm. This is especially important for near-term quantum computers that suffer from various sources of error that propagate with the circuit depth. Parametrized gates in particular have found use in both near-term algorithms and circuit compilation. The one- and two-qubit versions of these gates have been demonstrated on various computing architectures. The neutral atom platform has the capability to implement native $N$-qubit gates (for $N\geq 2$). However, one needs to first find the control functions that implement these gates on the hardware. We study the numerical optimization of neural networks towards obtaining families of controls – laser pulses to excite an atom to Rydberg states – that implement phase gates with one and two controls, the $\mathrm{C_{1}P}$ and $\mathrm{C_{2}P}$ gates respectively, on neutral atom hardware. The pulses we obtain have a duration significantly shorter than the loss time scale, set by decay from the Rydberg state. In addition, they do not require single-site addressability and are smooth. Hence, we expect our gates to have immediate benefits for quantum algorithms implemented on current neutral atom hardware.
@@ I Introduction
To implement these gates on hardware, however, one must devise pulses – that is, the time profile of the laser intensity and frequency – that can carry out the desired unitary transformation on the electronic states of the atoms. On the theoretical side, such pulses have been proposed for $k$-qubit Toffoli and fan-out gates (for $k>2$[13, 14, 15, 16, 6, 17], as well as for $k$-controlled $\mathrm{SWAP}$ gates [18, 19] and more general unitaries [20]. Such gates have also been demonstrated experimentally on the Rydberg platform [15, 21, 22]. Notably, Cao et al. implemented $N$-qubit rotation gates (for $N$ up to 9) in order to generate cascaded GHZ states [21] by placing $N$ atoms within interacting range of each other. The pulses devised to implement this gate are also time-optimal – meaning they are the shortest possible pulses that implement the corresponding unitary. Such pulses that implement the gate directly are significantly shorter in duration than implementing the decomposition of the gate in terms of one- and two-qubit gates [23], and are obtained numerically using a class of methods termed quantum optimal control (QOC) [13, 24].
@@
In particular, we highlight the variational neural network (NN) methods that have gained prominence in quantum control problems [25, 26, 27, 28, 29, 23]. These networks are variational ansatze with a structure inspired by the networks of neurons used for learning in the brain. Importantly, they are universal approximators of multivariate, continuous functions [30], and the exact gradients over their weights can be computed numerically [31]. This means that NN-based methods can generalize to challenging QOC tasks, including control of driven systems [29] and systems including quantum feedback [26].
Recent studies have also investigated the parametrized versions of multiqubit gates – gates that implement some unitary $U(\bm{\phi})$, for $M$ parameters $\bm{\phi}=(\phi_{1},\phi_{2},\dots,\phi_{M})$. In terms of variational algorithms, there are proposed applications of such gates towards simulating lattice gauge theories [32] and combinatorial optimization [33]. Further, for a given circuit, the inclusion of $k$-controlled phase, or $\mathrm{C}_{k}\mathrm{P}$ gates (for $k\geq$ 2) in the gate set can reduce the gate count and the total circuit evaluation time of the compiled circuit [12], thus reducing infidelities – as most error sources propagate in time [13]. This reduction is especially significant for the current era of quantum hardware, termed the noisy intermediate-scale quantum (NISQ) era. In these devices, the coherence times of the qubits are limited, and there are multiple sources of gate infidelity.
We note that preliminary work exists on $\mathrm{C}_{k}\mathrm{P}$ protocols for neutral atoms, in particular for the $\mathrm{C_{1}P}$ [34, 15] and $\mathrm{C_{2}P}$ [35] gates. A common limitation across these studies is the necessity of re-optimizing control parameters for each specific gate angle $\phi$. Such protocols return control function(s) that implement the gate at a fixed $\phi$ – whereas the complete protocol must be a family of controls, that is, continuous (or piecewise continuous) in $\phi$ [23]. We will discuss these studies further in Sec. II.1.
We investigate the problem of finding families of pulses that implement the $\mathrm{C_{1}P}$ and $\mathrm{C_{2}P}$ gates on Rydberg hardware. Employing numerical techniques from QOC, we find pulses for both gates that are fast – more specifically, the $\mathrm{C_{1}P}$ pulse we obtain is time-optimal, and the $\mathrm{C_{2}P}$ pulse operates on a time scale comparable to the time-optimal solution. These pulses are also smooth, and require only global addressing – the same pulse is sent to all atoms – and hence can be implemented on NISQ devices.
To the best of our knowledge, our paper hence presents the first $\mathrm{C}_{k}\mathrm{P}$ pulse families on the Rydberg platform. Further, the techniques we use for the optimization can be extended to gates over more control qubits, as well as to fast quantum state preparation for parametrized states.
@@ II.1 Multiqubit gates
Multiqubit gates, including $\mathrm{C_{2}P}$ gates, have been realized natively on various platforms. In Appendix C we present an overview of platform-specific methods towards generating these gates. One can always implement such transformations by decomposing them into a product of one- and two-qubit gates available on the hardware – this is visualized for the $\mathrm{C_{1}P}$ gate in Fig. 2. However, this approach has a significantly longer gate duration [23, 39] than the control-level approach we adopt, and we also present an in-depth comparison of these two approaches for our problem in Appendix C.
We note that as $\mathrm{P(\pi)=Z}$, $\mathrm{C}_{k}\mathrm{P}(\pi)$ is equivalent to a $\mathrm{C}_{k}\mathrm{Z}$ gate. Hence, the problem we consider is the generalized version of the problem of obtaining a $\mathrm{C}_{k}\mathrm{Z}$ pulse [13, 40, 41, 34, 42].
@@
and, defining the operator
$\hat{O}_{k}^{i,j}=\begin{cases}|r\rangle\langle r|,&\text{if }k\in\{i,j\},\\ \mathds{1},&\text{otherwise.}\end{cases}$ (10)
the interaction term is
$H_{\mathrm{int}}^{i,j}=-\frac{C_{6}}{|\bm{R}_{i}-\bm{R}_{j}|^{6}}\bigotimes_{k=1}^{N}\hat{O}_{k}^{i,j}.$ (11)
@@
For an $N$-qubit system in the blockade regime, evolving under the Hamiltonian from Eq. (8), one can realize unitaries that map each computational basis state to itself up to a phase [34]. This means that the system is well suited to realizing $\mathrm{C}_{N-1}\mathrm{P}$ gates [12] – upto a global single qubit $\mathrm{R_{Z}}$ rotation. The dynamics in the blockade regime can be understood in terms of two-level systems (TLS): two and three for the $\mathrm{C_{1}P}$ and $\mathrm{C_{2}P}$ gate, respectively. We illustrate these systems in Fig. 1. The TLS are closed systems – independent of each other, they partition the Hilbert space into subspaces with populations that do not mix [34].
@@
We note, however, that the blockade strength is finite in the experiment [as computed in Eq. (11)], and this is a potential source of infidelity for gates optimized at $B\to\infty$. For the $\mathrm{C_{1}P}$ gate, we will hence simulate the dynamics (during the optimization) under the full Hamiltonian from Eq. (8) with $B=21.1$, an experimentally realistic value [40, 15] – as opposed to simulating $H_{\mathrm{eff}}$. This is done in order to mitigate the finite blockade errors [41].
@@
When reporting the final fidelities in Sec. III, one must also consider the losses arising from spontaneous emission and blackbody radiation (BBR) from the Rydberg state [45] at $300\mathrm{K}$. To simulate this effect, we add a non-Hermitian term [40, 41]
$H_{\mathrm{decay}}=i\frac{\Gamma_{n=61}}{2}\ket{r}\bra{r},$ (15)
to $H_{1q}$ in Eq. (7), which assumes that all decay occurs to states outside of the computational subspace – an assumption justified for the system we studied in this article [40, 46]. Here $\Gamma_{n=61}$ is chosen as the decay width of the $5\mathrm{s}61\mathrm{s}\,^{3}\mathrm{S}_{1}$ state in Sr, corresponding to a lifetime $1/\Gamma_{n=61}=96.5\mathrm{\mu s}$ [41]. This term is subsequently used to construct the full Hamiltonian $H_{Nq,\Gamma}$, in a manner similar to Eq. (8).
@@ II.2 Numerical optimization
Sauvage et al. investigated the generation of parametrized gates by training neural networks – used as variational ansatze – to model families of pulses [23]. We build upon their approach, and adapt it to our neutral atom system. As we will motivate in this section, compared to GRAPE, the NN only has to be trained once, and then outputs a high-fidelity pulse for any angle in the domain of the gate without requiring further optimizations.
@@
for $\mathcal{T}$ the time-ordering operator. Defining the projection operator $P$ onto the computational subspace [49],
$P=\sum_{i_{1},i_{2},\dots i_{N}\in\{0,1\}}\ket{i_{1}i_{2}\dots i_{N}}\bra{i_{1}i_{2}\dots i_{N}},$ (17)
the cost function at each iteration is
$J=1-\frac{1}{4^{(k+1)}}\braket{}{U_{\mathrm{C}_{k}\mathrm{P}}(\phi)^{{\dagger}}(PU_{\mathrm{out},k}(\phi)P)}{{}^{2}}_{\phi},$ (18)
where ||.||||.||| | . | | is the Hilbert-Schmidt norm and .$\braket{.}_{\phi}$ implies we compute the mean over samples of $\phi$ drawn randomly from $\mathrm{dom}(\mathrm{C}_{k}\mathrm{P})$ in each iteration. We will refer to $J=\braket{(1-F)}_{\phi}$ as the mean infidelity over the domain of the protocol. We then train the weights of the NN to minimize the cost function.
@@ III Results
As discussed in Sec. II.1, to compute the infidelity of our pulses in a realistic experimental setup, we simulate decay from the Rydberg state, with estimated lifetime $\tau=96.5~{}\mathrm{\mu s}(\simeq 6\times 10^{3}/\Omega_{\mathrm{max}})$ for the $n=61$ state [41]. Unless specified otherwise, a realistic blockade strength of $B=21.1$ [40] is also included in the simulation. We consider a Rabi frequency of $\Omega_{\mathrm{max}}=2\pi\times 10\mathrm{MHz}$ for Sr, which is within reach in the near term [3, 40]. Calculations of infidelities arising from the decay of |$\ket{r}$, and from finite blockade, are discussed in Appendix B. The total mean infidelity $\braket{(1-F)}_{\phi}$ was computed over 200 samples randomly drawn from $\mathrm{dom}(\mathrm{C}_{k}\mathrm{P})$.
@@ III.1 $\mathrm{C_{1}P}$ gate
Simulating the errors as discussed above, we obtain a infidelity $\braket{(1-F)}_{\phi}=3.4\times 10^{-4}$ for the two-qubit $\mathrm{C_{1}P}$ gate over $\mathrm{dom(C_{1}P)}$ – with the main error arising due to Rydberg state decay (see inset of Fig. 4(b)). We also report the gate infidelity averaged over states in the Haar measure [57], $\braket{(1-F^{\prime})}_{\phi,\psi}=2.7\times 10^{-4}$.
@@ III.2 $\mathrm{C_{2}P}$ gate
We hence study the strategy of using the $\mathrm{C_{2}P}$ NNs we trained at $B\to\infty$, as initial guesses for $\mathrm{C_{2}P}$ NNs trained at $B=21.1$, and then re-run the optimization as discussed in Sec. II.2. The pulses we obtain from this optimization reduce the average finite blockade infidelity to $\braket{(1-F)_{\mathrm{int}}}_{\phi}=4.0\times 10^{-4}$, and the total infidelity $\braket{(1-F)}_{\phi}=1.45\times 10^{-3}$ – where the decay from the Rydberg state is the main source of infidelity. The fidelity averaged over the states in the Haar measure, $\braket{(1-F^{\prime})}_{\phi,\psi}=1.28\times 10^{-3}$. The results of this optimization are visualized in Fig. 6, in Appendix B.
@@ IV Conclusion
As discussed in the introduction, gate sets containing $\mathrm{C}_{k}\mathrm{P}$ gates can reduce circuit depth in quantum compilation, when compared to standard gate sets [12]. Our results are of interest for this area of research, as they enable the realization of the $\mathrm{C}_{k}\mathrm{P}$-compiled circuits on neutral atom hardware. We are currently investigating resource-efficient state preparation schemes using these gates.
@@ Appendix B Losses from finite blockade and Rydberg state decay
for $F,F_{\mathrm{decay}}$ the fidelities are obtained by simulating the pulses under $H_{Nq}$ [see Eq. (8)], and $H_{Nq,\Gamma}$ respectively. The infidelity arising from finite blockade was defined for three qubits, in a manner similar to Eq. (24), $(1-F)_{\mathrm{int}}=F_{\infty}-F_{\mathrm{fin}}$, for $F_{\infty},F_{\mathrm{fin}}$ the fidelities of the pulses simulated for $H_{\mathrm{eff}}$ and $H_{3q}$ respectively. This infidelity is visualized as a function of $\phi$ in Fig. 6.
@@
Appendix C Native gates and platforms
@@
It is also worth highlighting the native $N$-qubit gate protocols (for $N>2$) that have been proposed or demonstrated on various quantum platforms. This includes native implementations of $N$-qubit Toffoli gates [69] and Mølmer-Sørensen entanglement schemes [70] on trapped ion systems. Further, $\mathrm{C_{2}P}$ protocols have also been implemented on the superconducting platform [71].
In terms of gates, it is straightforward to devise native $N$-qubit protocols on the neutral atom platform. This is in part due to the long range interactions, and partly due to reconfigurable arrays–the ability to move the atoms around the tweezer during computation, allowing for gates over distant qubits without additional lossy SWAP operations [5]. As noted in the Introduction, gates on up to $9$ qubits have been experimentally demonstrated [21], and several protocols have been proposed for $\mathrm{C}_{k}\mathrm{Z}$ gates with an arbitrary number of controls $k$.
Trapped ion systems enjoy similar benefits to multiqubit gate implementations. In contrast, such gates on superconducting quantum computers currently require significant modifications to the hardware and are difficult to scale to multiple control qubits [71]. Recent studies also demonstrate native multiqubit gates with quantum dots [72] and solid-state systems [73], and hence a systematic comparison of these platforms is a timely question, made challenging by their rapid pace of development.